OpenStudy (anonymous):

Vilfredo Pareto (1848-1923) observed that most of the wealth of a country is owned by a few members of the population. Pareto's Principle is log P = log c − k log W where W is the wealth level (how much money a person has) and P is the number of people in the population having that much money. (a) Solve the equation for P.

OpenStudy (anonymous):

log(p) = log(c/w^k) p = c/(w^k)

OpenStudy (anonymous):